How long will Php60,000 accumulate to Php62,430 if the simple interest rate agreed is 1.50%?

1
Expert's answer
2021-11-21T15:10:31-0500

"t =(\frac{1}{r}) (\frac{a}{p}-1)"

T= time in years

R= rate of interest = 1.50% = 0.015

A= final amount = 62,430

P= initial amount =60,000


"t = (\frac{1}{0.015}) (\frac{62430}{60000}-1)"


"t=(66.6667)(1.0405-1)"


"t=(66.6667)(0.0405)"


"t= 2.7"

It will take approximately 3 years